### v1

# Cluster-Enhanced Federated Graph Neural Network for Recommendation

Abstract
Personal interaction data can be effectively modeled as individual graphs for each user in recommender systems. Graph Neural Networks (GNNs)-based recommendation techniques have become extremely popular since they can capture high-order collaborative signals between users and items by aggregating the individual graph into a global interactive graph. However, this centralized approach inherently poses a threat to user privacy and security. Recently, federated GNN-based recommendation techniques have emerged as a promising solution to mitigate privacy concerns. Nevertheless, current implementations either limit on-device training to an unaccompanied individual graphs or necessitate reliance on an extra third-party server to touch other individual graphs, which also increases the risk of privacy leakage. To address this challenge, we propose a Cluster-enhanced Federated Graph Neural Network framework for Recommendation, named CFedGR, which introduces high-order collaborative signals to augment individual graphs in a privacy preserving manner. Specifically, the server clusters the pretrained user representations to identify high-order collaborative signals. In addition, two efficient strategies are devised to reduce communication between devices and the server. Extensive experiments on three benchmark datasets validate the effectiveness of our proposed methods.
Index Terms:
Federated Learning, Graph Neural Network, Recommender System, Privacy-Preserving, Cluster
I Introduction
In recommendation systems, the interaction data for each user can be effectively represented as a individual graph structures [1]. These personalized graphs capture the unique relationships and interactions between the user and various items, allowing for a more nuanced and accurate modeling of user preferences[2]. The advanced capability of Graph Neural Networks (GNNs) to aggregate individual user interaction graphs into a comprehensive global interactive graph has led to their widespread application in recommender systems [3]. However, this centralized learning methodology necessitates the aggregation and centralized storage of substantial volumes of users’ confidential data[4]. Despite its advantages, this method has several limitations: 1)The data involved is highly sensitive and the transmission of users’ private interaction data beyond local devices poses a significant risk of privacy leakage[5]. 2)Centralized data storage necessitates that third-party devices possess sufficient storage capacity. 3)Stringent data protection regulations, such as the General Data Protection Regulation (GDPR)[6] have been enacted, prohibiting commercial entities from collecting, exchanging, or selling user data without explicit user consent. These limitations make it difficult to deploy centralized learning methods in real-life scenarios.
Fortunately, the Federated Learning (FL) framework proposed by Google[7] addresses these issues. FL allows multiple users to collaboratively train a global model, with users needing to upload only model parameters instead of raw data. In such scenarios, users are confined to constructing first-order subgraphs from their own interaction data, precluding the exploitation of information from other users with similar interests, specifically higher-order collaborative signals. This constraint prevents GNNs from capturing high-order information, leading to degraded recommendation performance[8]. To address this issue, algorithms like FedGNN[9] have been proposed, but it brought about risks of privacy leakage and increased communication burdens. Acquiring advanced collaborative signals while maintaining user privacy has emerged as a critical challenge.
In light of the aforementioned challenges, we propose a Cluster-enhanced Federated Graph Neural Network framework for Recommendation(CFedGR). Within this framework, user features extracted by GNNs are aggregated into clusters. Given that these features integrate information from interacted items, users within the same cluster demonstrate analogous preferences, effectively serving as each other’s collaborative context. To alleviate the communication overhead between server and local devices during the model training, we devise a sampling strategy to select representative users from each cluster for training, and restrict the number of collaborative neighbors to the top k most similar ones. To sum up, our contributions are summarized as follows:
  1. 1.
    We propose a cluster-enhanced federated graph neural network framework for recommendation, which effectively captures users’ high-order collaborative information without compromising user privacy.
  2. 2.
    We propose a clustering technique to augment users’ local first-order interaction subgraphs and devise two straightforward yet effective strategies to alleviate communication overhead between server and devices.
  3. 3.
    Extensive experiments, conducted on three benchmark datasets used in recommendation systems, have validated the effectiveness of our proposed methods.
II RELATED WORKS
II-A Graph Neural Network
GNN have garnered extensive research attention in recent years owing to their exceptional capability to efficiently model graph data, which has prompted their wide adoption in different graph-based tasks[10], such as node classification[11], link prediction[12], and graph classification[13]. By employing the technique of message passing defined as follows:
$\mathbf{h}_{v}^{(k+1)}=\sigma\left(\mathbf{W}\cdot\text{AGGREGATE}\left(\{\mathbf{h}_{u}^{(k)}:u\in\mathcal{N}(v)\}\right)\right)$ (1)
where $\mathbf{h}_{v}^{(k+1)}$ represents the feature vector of node $v$ at layer $k+1$, and $\mathbf{h}_{u}^{(k)}$ is the feature vector of node $u$ at layer $k$. The set $\mathcal{N}(v)$ consists of the neighbor nodes of node $v$. The AGGREGATE function collects information from the neighbors, utilizing methods such as summation, averaging, or taking the maximum. $\mathbf{W}$ is a matrix of learnable weights and $\sigma$ signifies the activation function. GNN can iteratively combine the features of neighboring nodes with its own features to update the node’s representation[14]. For example, Kipf et al.[15] proposed a scalable approach for semi-supervised learning on graph-structured data using Graph Convolutional Networks (GCNs). It employs a localized first-order approximation of spectral graph convolutions to efficiently capture local graph structure and node features. Hamilton et al.[16] introduced GraphSAGE, which enhances GNN by sampling and aggregating features from a node’s local neighborhood. Veličković et al. [17] presented Graph Attention Networks (GAT) that employ an attention mechanism to assign different importances to nodes in a neighborhood, allowing the model to focus on the most relevant parts of the graph. He et al.[18] proposed LightGCN, by eliminating feature transformation and nonlinear activation, they simplify traditional GCNs, creating a more efficient model that focuses solely on neighborhood aggregation.
II-B Federated Recommendation Systems
Federated Learning (FL) allows multiple clients to collaboratively train machine learning models while safeguarding data privacy and has seen a significant rise in application to recommendation systems in recent years, which is known as Federated Recommender Systems (FedRecs) [19]. In this paradigm, rather than transmitting their private data, clients engage in the training of local model using data stored locally and subsequently transmit their model parameters or gradients to a central server, which then aggregates this information to derive a comprehensive global model, effectively mitigating the risk of data leakage. FedRecs can generally be categorized into two approaches: federated recommendation based on matrix factorization and federated recommendation based on graph neural networks. Matrix factorization is a technique commonly used in recommendation systems. It works by decomposing the user-item rating matrix into two lower-dimensional matrices, the user matrix U and the item matrix V, such that their product approximates the original rating matrix. For example, Federated Collaborative Filtering (FCF) [20] suggested each client utilizes local data and item embeddings retrieved from the server to update user embeddings, ensuring that user data never leaves the client device. Du,Yongjie et al.[21] introduced a randomized response mechanism to further enhance privacy protection. Recommendation methods based on graph neural networks leverage these networks to derive latent representations of users and items, which are consequently employed for downstream recommendation tasks. For instance, FedSoG[22] adopts a relational attention mechanism to distinguish the contributions of neighboring users and interacted items to the user’s embedding, effectively addressing heterogeneity in the process. A similar work to ours is FedGNN [9]. It utilizes a trusted third-party server and homomorphic encryption to identify users’ neighbors and augment the local subgraph. However, this approach still carries the risk of leaking user privacy and requires higher communication overhead.
III PROPOSED METHOD
Figure 1: Framework of the CFedGR
III-A Problem Formulation
Consider $U=\{u_{1},u_{2},\ldots,u_{M}\}$ and $T=\{t_{1},t_{2},\ldots,t_{N}\}$ as the sets of users and items, with $M$ and $N$ representing the number of users and items, respectively. The embeddings of users and items are denoted by $\mathbf{E}_{u}\in\mathbb{R}^{M\times d}$ and $\mathbf{E}_{i}\in\mathbb{R}^{N\times d}$, where $d$ indicates the dimensionality of these vectors. Denote the neighbors of user $u_{i}$ as $N_{i}=\{n_{i,1},n_{i,2},\ldots,n_{i,p}\}$, where $p$ represents the number of neighboring users. Users assign ratings to the items they interact with. For example, the rating given by user $i$ to item $j$ is denoted as $r_{ij}$. For each user $u_{i}$, let $t_{i}=\{t_{i,1},t_{i,2},\ldots,t_{i,q}\}$ represent the set of items with which the user has interacted, with the corresponding ratings represented as $r_{i}=\{r_{i1},r_{i2},\ldots,r_{iq}\}$, where $q$ denotes the number of items the user has rated. Based on the interaction data from all users, a rating matrix $\mathbf{R}\in\mathbb{R}^{M\times N}$can be obtained. To mitigate the risk of data leakage, users’ interaction data is retained locally on their devices. Accordingly, there exists solely a first-order user-item graph, consisting solely of each user and the items they have interacted with. Our objective is to identify each user’s collaborative neighboring users without compromising user privacy and utilize the neighbors’ information to predict the ratings($r\in\mathbf{R}\setminus\sum_{i\in U}r_{i}$) for items that the users have not previously interacted with .
III-B Overview of the Framework
This section introduces the proposed framework, as illustrated in the Fig. 1. The framework consists of two main components: the client-side and the server-side. The client-side is responsible for training a local GNN with data stored on local devices. We represent the embeddings of the user, the user’s neighbors, the items interacted with and the negative items by the user as $e_{u_{i}}$, $\{e_{u_{i,1}},e_{u_{i,2}},\ldots,e_{u_{i,p}}\}$, $\{e_{t_{i,1}},e_{t_{i,2}},\ldots,e_{t_{i,q}}\}$, and$\{e_{n_{i,1}},e_{n_{i,2}},\ldots,e_{n_{i,s}}\}$, respectively. Subsequently clients upload both the embedding gradients and model gradients to the server. The server is tasked with three main responsibilities: determining the users eligible for the training, aggregating the gradients uploaded by the clients to update global model and embeddings, and discerning the collaborative neighbors associated with each user. Next, we will detail the coordination mechanism between the server-side and client-side.
At the outset of each training round, the server selects a subset of users for participation. An embedding layer is deployed on the server to map user IDs and item IDs into embeddings. Acknowledging the initial imprecision of user embeddings, T rounds of pre-training are performed without utilizing neighbor embeddings. Following this phase, the server undertakes clustering based on the user embeddings to discern the neighbors for each user. Upon obtaining the embeddings of users, collaborative neighbors, and interacted items, the client feeds these embeddings into the local GNN to derive more expressive embeddings. In the subsequent step, the predicted ratings are computed by taking the inner product of the user embeddings and item embeddings according to (2).
$\hat{r}_{ij}=\mathbf{u}_{i}\cdot\mathbf{v}_{j}$ (2)
After obtaining all the predicted ratings $(\hat{r}_{i1},\hat{r}_{i2},\ldots,\hat{r}_{iq})$ for user $u_{i}$, we use (3) to compute the loss with the original ratings $(r_{i1},r_{i2},\ldots,r_{iq})$.
$\text{Loss}=\frac{1}{q}\sum_{k=1}^{q}(r_{ik}-\hat{r}_{ik})^{2}$ (3)
$\theta$ is defined as the GNN parameters. Derived from the loss function, the gradients of GNN, user embeddings and item embeddings $\nabla_{\theta}L$, $\nabla_{E_{u}}L$, and $\nabla_{E_{i}}L$ are then uploaded to the server. Once the gradients are received from all clients, the server aggregates them using the following formula:
$\mathbf{\nabla\mathcal{L}}^{\prime}=\sum_{k=1}^{N}\frac{n_{k}}{n}\nabla\mathcal{L}_{k}$ (4)
where $n_{k}$ is the number of interaction data for user $k$, and $n$ is the total number of interaction data for all users, the $\nabla\mathcal{L}_{k}$ is the gradients of user $u_{k}$. We designate the learning rate as $l$. According to (5), (6), and (7), the updated GNN parameters $\theta^{\prime}$, user embeddings $\mathbf{E}_{u}^{\prime}$ and item embeddings $\mathbf{E}_{i}^{\prime}$ are obtained.
$\theta^{\prime}=\theta-l\nabla_{\theta}\mathcal{L}^{\prime}$ (5)
$\mathbf{E}_{u}^{\prime}=\mathbf{E}_{u}-l\nabla_{\mathbf{E}_{u}}\mathcal{L}^{\prime}$ (6)
$\mathbf{E}_{i}^{\prime}=\mathbf{E}_{i}-l\nabla_{\mathbf{E}_{i}}\mathcal{L}^{\prime}$ (7)
Finally the server broadcasts the updated GNN and embeddings to the clients for the next round of training. In the following sections, we provide a detailed explanation of several important modules within the framework.
III-C Local User-Item Graph Augmentation via Clustering
In the context of federated learning, each user exclusively retains data pertaining to the items they have interacted with, acquiring information about a user’s collaborative neighbors while ensuring their privacy remains intact is a considerable challenge. Under current methodologies such as FedGNN [9], which incorporate an additional third-party server alongside the central server and utilize homomorphic encryption, clients are able to upload encrypted user embeddings and item IDs. However, this approach not only escalates the deployment costs and network communication burden but also poses a significant privacy risk if the central and third-party servers collude. To address this issue, we propose a clustering-based method to expand the local user-item graph, which can reveal user’s neighbors while preserving user privacy.
Firstly, the server that holds all user embeddings performs clustering on these embeddings. To prevent performance degradation caused by inaccurate user embeddings at the initial stage, we conduct clustering after pre-training for $T$ rounds. Numerous clustering algorithms available for this purpose, including the K-means algorithm[23], hierarchical clustering[24], and the DBSCAN algorithm[25]. In our method, we employ the K-means clustering algorithm. After that, users with similar interests are grouped into the same cluster. Except for the user themselves, the remaining users in the cluster are considered the user’s collaborative neighbors. Proceeding with this, the IDs of other users in the cluster are sent to each user by server, and edges are added between the user and the received neighbor nodes. Throughout this process, the server does not receive any private data from the users, nor is there any data sharing between users, which ensures that the local first-order user-item graph can be expanded while maintaining user privacy.
III-D Strategies for Communication Efficiency
To further reduce the communication overhead between the server and devices, we design two simple yet effective methods. Firstly, instead of the random user selection commonly used in most federated recommendation systems, we cluster all user embeddings on the server-side and select a proportional number of users from each cluster to participate in the training. This ensures that the model is exposed to a wider variety of user types and behavior patterns during training, which aids in better learning and generalization of the model. Secondly, during the user neighbor matching phase, the server does not return the IDs of all users within the same cluster, as not all of them contribute to constructing the user’s preferences. Instead, server selects the top k users with the highest similarity to the user based on the similarity matrix, which helps to further reduce the communication between the server and the users.
III-E User Privacy Protection
In the model training phase, users need to upload gradients for interacted items. However, only items with which the user has interacted have non-zero item gradients. As a result, the server can potentially reconstruct the user’s interaction history based on these non-zero gradients. To safeguard against the disclosure of users’ interaction histories, we sample $n$ negative items from those not interacted with by the user and generate gradients for these negative samples using a Gaussian distribution. These generated gradients are then uploaded along with the actual item gradients, making it difficult for the server to distinguish the user’s true interaction history. Denoted the real item gradients as $\mathbf{g}_{\text{real}}$, with the mean and variance of these gradients represented by $\mathbb{E}[\mathbf{g}_{\text{real}}]$ and $\mathbb{V}[\mathbf{g}_{\text{real}}]$, respectively. $\mathbf{z}$ is a random vector following a normal distribution. According to (8),we can obtain the gradients for the negative samples.
$\mathbf{g}_{\text{neg}}=\mathbb{E}[\mathbf{g}_{\text{real}}]+\sqrt{\mathbb{V}[\mathbf{g}_{\text{real}}]}\cdot\mathbf{z}$ (8)
As the model gradients uploaded by users contain private information such as user preferences, they also pose a risk of privacy leakage. To mitigate this, we employ local differential privacy by adding noise to these gradients before uploading them. This ensures that the server cannot infer users’ private information from the gradients.
$g^{\prime}=\text{clip}(g,\delta)+\text{Laplace}(0,\lambda)$ (9)
where $\text{clip}(g,\delta)$ clips the gradient $g$ to ensure its norm does not exceed $\delta$, $\text{Laplace}(0,\lambda)$ is Laplace noise with mean 0 and scale $\lambda$.
IV EXPERIMENTS
IV-A Experimental Settings and Comparisons
In our experiments, three benchmark datasets in recommendation systems are selected to evaluate the recommendation performance: FilmTrust, MovieLens-100k and Douban. The statistical information of these datasets is shown in Table I.
TABLE I: Statistics of the Dataset
Dataset Users Items Ratings Sparsity
Filmtrust 1508 2071 35497 98.85%
MovieLens-100K 943 1682 100000 93.70%
Douban 3000 3000 136891 98.47%
Each dataset is split into training, validation, and test sets with a ratio of $70\%$, $10\%$, and $20\%$, respectively. The learning rate is optimized within the range of {0.01, 0.05, 0.1}, while the decay rate is explored within {0.0001, 0.0005, 0.001}. Each iteration engages 256 users, with embeddings mapped into 128 dimensions. We establish a gradient clipping threshold of 0.2 and set the noise scale to 0.1. Additionally, the $k$-value is set to 200 and 1000 negative samples are utilized. A two-layer Graph Attention Network (GAT) serves as the local GNN. Performance is assessed using two classic metrics: Mean Absolute Error (MAE) and Root Mean Square Error (RMSE), with lower values indicating better results. Baseline models from both centralized and federated learning frameworks are employed to evaluate recommendation efficacy.
Centralized Learning:
  • NCF[26]: It combines matrix factorization and multi-layer perceptron to learn the interaction function between users and items.
  • GraphRec[27]: It integrates graph feature information into matrix factorization.
  • Glocal-K[28]: This approach pre-trains a localized kernelized weight matrix and trains an autoencoder.Then fine-tunes the autoencoder using global convolutional kernels.
TABLE II: Performance comparison of different methods
Method Filmtrust MovieLens-100K Douban
MAE RMSE MAE RMSE MAE RMSE
NCF 0.6351 0.8174 0.7354 0.9321 0.5964 0.7540
GraphRec 0.6554 0.8561 0.7128 0.9033 0.5788 0.7379
GLocal-K 0.6146 0.8026 0.6933 0.8895 0.5826 0.7428
MetaMF 0.6901 0.8756 0.7941 0.9990 0.6537 0.8259
FedNCF 0.6854 0.8613 0.7890 0.9922 0.6470 0.8134
FedGNN 0.6313 0.8162 0.7705 0.9646 0.6477 0.8036
Ours 0.6351 0.8170 0.7725 0.9661 0.6484 0.8067
(a) MAE performance
(b) RMSE performance
Figure 2: Ablation study without collaborative neighbors
Federated Learning:
  • FedNCF[29]: It is a federated version of NCF, where user and item embeddings are shared on the server.
  • MetaMF[30]: It incorporates a collaborative memory module and a meta-recommendation module for producing personalized item embeddings and recommendation models.
  • FedGNN[9]: It utilizes a third-party server and homomorphic encryption to find neighboring users, which enhances the feature representation of users.
IV-B Analysis of Experimental Results
The experimental results on all datasets are shown in Table II, from which we can make the following observations:
  1. 1.
    Although centralized learning methods achieve the lowest error rates, they pose significant risks to user privacy and face considerable challenges in real-life deployment due to data collection difficulties. In contrast, our method only slightly compromises performance while ensuring user privacy, thus offering a superior trade-off between privacy protection and recommendation accuracy.
    TABLE III: Number of neighbors required by FedGNN
    Datasets of neighbors
    Filmtrust 1094
    Douban 815
    MovieLens-100K 860
  2. 2.
    In contrast to other federated recommendation models that do not incorporate high-order collaborative neighbor information, such as MetaMF and FedNCF, our model delivers superior prediction performance across all datasets. This demonstrates the effectiveness of our proposed framework and highlights the critical role of neighbor information in enhancing recommendation systems.
  3. 3.
    Despite FedGNN showing marginally better than our model, the difference is minimal, with a maximum improvement of only 0.3%. However, FedGNN necessitates an additional server, which not only heightens the risk of privacy leakage but also increases server deployment costs and communication burdens. Furthermore, as shown in Table III, the number of neighbors that FedGNN needs to transmit is significantly higher than in our model. Our method reduces neighbor communication by approximately 4 to 5 times, achieving a superior balance between communication overhead and recommendation performance.
IV-C Hyperparameter Analysis
TABLE IV: Performance comparison on different values of N
of clusters MovieLens-100K Douban
MAE RMSE MAE RMSE
N=5 0.7822 0.9771 0.6846 0.8476
N=10 0.7826 0.9779 0.6839 0.8456
N=20 0.7815 0.9767 0.6831 0.8445
(a) MovieLens-100K
(b) Douban
Figure 3: Performance comparison on different values of K
In this section, we discuss the impacts of hyperparameters, specifically the number of clusters $N$ and the value of top $K$. We conducted experiments on two relatively large datasets, MovieLens-100K and Douban. Table IV demonstrates that performance remains stable across various cluster numbers when the top $K$ is fixed at 40, which alleviates the burden of parameter tuning. Furthermore, as illustrated in Fig. 3, RMSE gradually decreases with increasing $K$. However, once $K$ reaches 300, the reduction in error becomes negligible. To minimize communication overhead, we recommend setting the $K$-value to 200.
IV-D Ablation Study
To verify the effectiveness of the clustering algorithm, we conducted experiments under identical settings without its use. The results, presented in Fig. 2, show that performance is significantly worse without the clustering algorithm. This demonstrates the effectiveness of our clustering algorithm and underscores the importance of high-order collaborative neighbor information in recommendation systems.
V Conclusion
In this paper, we emphasize the significance of high-order collaborative information in federated recommendation systems and propose a cluster-enhanced federated graph neural network framework for secure item recommecdation. Our proposed framework clusters users with similar preferences together to derive high-order collaborative neighbors while preserving user privacy. To alleviate communication burden, we designed a sampling strategy that selects representative users from each cluster for each training round. Additionally, the server captures only the top k neighbors with the highest similarity to the user within the cluster, instead of considering all neighbors. Comprehensive experiments on three datasets confirm that our framework achieves an exemplary balance between privacy protection and recommendation performance.